OpenStudy (anonymous):

Solve using the multiplication principle. -6a = 42

OpenStudy (anonymous):

Multiply both sides by \(\large -\frac{1}{6}\).

OpenStudy (anonymous):

But whats the answer?

OpenStudy (anonymous):

a= -7

OpenStudy (anonymous):

\[\cancel{-6}a\cdot \cancel{\frac{-1}{6}}=42\cdot (-\cancel{\frac{1}{6})} \implies a=\frac{-42}{6}=-7.\]

OpenStudy (anonymous):

Lol, there's no need for that last cancellation line.